OpenStudy (anonymous):

PROBLEM SOLVING USING SYSTEMS OF EQUATIONS Suppose your class sells gift wrap for $4 per package and greeting cards for $10 per package. You class sells 205 packages in all and receives $1084. Find the # of packages of gift wrap and the number of packages of greeting cards sold.

OpenStudy (lgbasallote):

let x = no. of gift wrap sold y = greeting card sold x + y = 205 4x + 10y = 1084 multiply the first equation by -10 (because it's easier lol) -10x - 10y = -2050 4x + 10y = 1084 -6x = -966 x = 161 y = 44 161 gift wraps were sold and 44 greeting cards were sold.

OpenStudy (anonymous):

thankyou so much, i only have 1 more problem!